OpenStudy (anonymous):

simplify by removing factors of 1 q^2-25/(q+5)^2

OpenStudy (anonymous):

\[ \frac{q^2-25}{(q+5)^2} = \frac{-5+q}{5+q} \]

OpenStudy (anonymous):

is that simplified form?

OpenStudy (aroub):

(q-5)(q+5) --------- (q+5)(q+5) q-5 ---- q+5 \[q \neq 5 \] & \[q \neq-5 \] sorryy!!